# Learning in Structured Stackelberg Games

### v1

@@ Abstract
We study structured Stackelberg games, in which both players (the leader and the follower) observe information about the state of the world at time of play. Importantly, this information may contain information about the follower, which the leader may use when deciding her strategy. Under this setting, we show that no-regret learning is possible if and only if the set of mappings from contexts to follower types that the leader uses to learn is not “too complex”. Specifically, we find that standard learning theoretic measures of complexity do not characterize learnability in our setting and we give a new dimension which does, which we term the Stackelberg-Littlestone dimension. In the distributional setting, we give analogous results by showing that standard complexity measures do not characterize the sample complexity of learning, but a new dimension called the Stackelberg-Natarajan dimension does. We then show that an appropriate empirical risk minimization procedure achieves the corresponding sample complexity.
@@ 1 Introduction
Many strategic interactions between a principal and an agent can be cast as Stackelberg games [36]. In a Stackelberg game, the principal (also called the leader) commits to a possibly randomized strategy, before the agent (also called the follower) observes this strategy and best-responds to it. Due to their effectiveness in modeling such leader-follower interactions, Stackelberg games have had significant real-world impact over the past decade, particularly in security domains [35]. For example, in protecting endangered species [38], a wildlife agency determines ranger patrol locations before poachers select targets. Coast guards decide on shoreline patrol routes in advance of illegal fishers’ actions [12] and, similarly, airport security [32] chooses checkpoint screening strategies before smugglers decide which entry points to use.
@@
In many of these real-world settings, the state of the world at time of play can greatly influence the utilities of the players and subsequently the outcome of the game. For example, defending some areas of the airport might be deemed of higher- or lower-importance based on the schedule of flights and the number of passengers on board. Similarly, weather patterns might make it harder to reach animals in certain parts of the park, while breeding seasons might make others more vulnerable thereby affecting their risk levels as targets. When contextual information becomes available over time, using it could help the leader design more effective strategies against the follower. This raises a natural question: given that the state of the world might change in current and future rounds of play, how should the leader in a Stackelberg game design a policy that adapts to new contexts coming in over time?
@@
One way is to study repeated Stackelberg games with side information [24], where at each interaction both the leader and the follower get to observe a context that affects their utilities. As in Balcan et al. [5], in this setting each follower that the leader faces might have different preferences and objectives, i.e. there might be different follower types. Harris et al. [24] study this problem and show that as long as either the followers or the contexts are independently drawn from a fixed and unknown distribution, no-Stackelberg-regret learning is possible. That is, the leader is guaranteed to achieve nearly as much utility as the best fixed policy in hindsight. In the fully adversarial regime, however, regret can grow linearly with the time horizon. Interestingly, this lower bound follows from the adversary’s ability to encode an arbitrarily complex relationship between the sequence of contexts and the follower types.
@@
In this work, we take a more granular view of this impossibility result and ask whether introducing additional structure between the contextual information and the follower types can enhance the leader’s ability to maximize returns, even when both are chosen adversarially to the leader. Namely, there could be correlation between the two, making the context predictive of the type of follower about to appear next. In wildlife protection for example, specific tire tracks might signal the presence of a particular poaching group and market demand trends might be indicative of which animals’ poachers are now more likely to appear. Similarly, in airport security, current affairs often raise or lower the risk of an attack on certain international flights.
Algorithms proposed by Harris et al. [24] and later work [9] are designed to handle changes in utility caused by the context, but crucially they assume no structure in how contexts relate to follower types. If, however, there is a fixed and unknown mapping between contexts and follower types, can the leader learn to exploit it to achieve sublinear regret? In this case, the leader m ust adapt her policy based on which follower types are more likely to appear, given the context and past observations. This poses an additional and possibly more complex learning task for the leader, which we study in this work.
@@ 1.1 Our Contributions
In this paper, we revisit the framework of Harris et al. [24] and take a learning theoretic approach to analyzing the leader’s task in predicting the follower type from contexts. Importantly, we find that existing notions of complexity do not capture the difficulty of learning in this setting—we obtain (tight) sample complexity and Stackelberg regret guarantees that are parameterized by new notions of complexity and that directly exploit the structure of Stackelberg games. Concretely, our main contributions in this work are as follows:
  1. 1.
@@
    We formalize online learning in Stackelberg games with side information, under the structural assumption that there exists a fixed, unknown mapping between contexts and follower types. In particular, we study how the leader, given access to a set of possible such mappings, can leverage it to learn an optimal policy. In particular, we study this problem under two different settings: (1) the realizable setting, in which there exists an optimal predictor of the followers’ types that the leader competes with; and (2) the agnostic setting, in which no such assumption is made.
  2. 2.
    We circumvent the impossibility result shown in Harris et al. [24] by providing algorithms with no-regret guarantees under the above structural assumption. We find that our problem is closely related to the problem of multiclass prediction, yet, interestingly, we show that our setting can be easier to learn in. In particular, we show that the (multiclass) Littlestone dimension, which quantifies the difficulty of learning in online multiclass prediction, can give loose mistake bounds in our setting. Motivated by this observation, we introduce a new notion of complexity, the Stackelberg-Littlestone dimension. We show that it characterizes learnability in our setting, and prove it to be at most as large as the multiclass Littlestone dimension. In fact, we give an example in which the latter is infinite, yet the former is zero. We also provide no-regret algorithms both for both the realizable and agnostic settings.
  3. 3.
    We introduce and formalize the study of distributional learning in Stackelberg games with side information. When the leader has access to a set of (context, follower type) examples, we show generalization with PAC-style bounds. Here too, we find that existing measures which characterize the complexity of multiclass prediction, such as the Natarajan dimension, do not characterize the sample complexity of our problem. Analogous to the online setting, we introduce the notion of Stackelberg-Natarajan dimension, with which we achieve an upper bound that is tight up to logarithmic factors.
@@ 1.2 Related Work
Stackelberg games. In the one-shot version of a Stackelberg game, Conitzer and Sandholm [16] prove that the problem of computing a Stackelberg equilibrium is NP-hard and they give an algorithm for maximizing the leader’s payoff. Other works focus on relaxing the assumption that all of the game’s parameters are known (e.g. utilities, number of actions) and aim to learn them instead [2, 31].
Our work is closely related to Bayesian Stackelberg games, in which the leader is playing against a single unknown follower type that is drawn from a given set of follower types [16]. This has been studied both in one-shot games, in which the leader must find a single optimal-mixed strategy [16, 30, 37] and in the repeated setting, in which the leader faces the same follower type sequentially and must learn from previous interactions [28, 26].
@@
Learning in Stackelberg games. The problem of learning an optimal strategy for the leader with a single unknown follower type has been studied by Letchford et al. [26] and Blum et al. [11]. These results have been extended by Peng et al. [31], who give faster rates and almost-matching lower bounds. Recently, Bacchiocchi et al. [2] relax some structural assumptions from previous work, including tie-breaking rules for the follower, and revise worst-case sample complexity results of learning the leader’s optimal mixed strategy. Recent research has explored many domains through the framework of Stackelberg games, including strategic classification, where classifiers account for individuals who strategically modify their features [23]; Bayesian persuasion, which models how a leader strategically reveals information to influence a follower’s decisions [25]; interactions with calibrated followers, who update their strategies based on past observations rather than making immediate best responses [21]; and settings with non-myopic followers, who optimize their actions over multiple time steps rather than reacting myopically [20, 15].
@@
Part of our work focuses on learning in online Stackelberg games. Online Stackelberg games with multiple follower types were introduced in Balcan et al. [5]. The authors show that when a series of unknown attacker types is chosen adversarially, playing the well-known Hedge algorithm  [19] over a special set of mixed strategies yields a no-regret algorithm.
@@
Our departing point is the work of Harris et al. [24], in which the authors study the problem of online learning in Stackelberg games with side information. The authors show that when either the sequence of follower types or the sequence of contexts are drawn independently from some fixed, unknown distribution, no-regret learning is indeed possible. However, when both of them are chosen adversarially, regret can scale linearly in the worst case, as the problem can be reduced from online learning linear thresholds. On the other hand, we make a structural assumption about the relationship between contexts and follower types: namely that it belongs to a class of functions that the leader has access to. With this, we show that learning is possible, and our guarantees depend on an appropriate measure of complexity of this function class.
@@
Multiclass learnability. Our work can also be viewed through the lens of multiclass learnability, as it suffices for the leader to learn the hypothesis class of the function mapping the contexts to a (finite) set of follower types. Then the leader can compute the optimal strategy against the predicted follower. Multiclass prediction has been studied extensively both in the ditributional ([29, 17, 13]) and in the online setting ([17, 22]). Indeed we adopt several tools from online multiclass prediction in our work. Interestingly, however, we show that learning in Stackelberg games with contexts is at most as hard as multiclass prediction, and in some cases, can be much easier. For example, in the online setting, we give an example in which the leader can learn to play optimally, under a hypothesis class that has infinite Littlestone dimension.
@@ 2 Problem Setting
We use $[N]:=\{1,2,\ldots,N\}$ to denote the set of natural numbers up to and including $N$. We use $\Delta(\pazocal{A})$ to denote the probability simplex over a (finite) set $\pazocal{A}$.
Stackelberg games with side information.
Consider a game between a leader and a follower. Before the game starts, some contextual information $\mathbf{z}\in\pazocal{Z}\subseteq\mathbb{R}^{d}$ is revealed to both players. The leader plays first and commits to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$, where $\pazocal{A}$ is the set of finite actions available to the leader. The follower best responds to the leader’s strategy, $\mathbf{x}$, and the context, $\mathbf{z}$, by playing some action $a_{f}\in\pazocal{A}_{f}$, where $\pazocal{A}_{f}$ is the finite set of follower actions. [^2] Note that without loss of generality the follower’s best response will always be a pure strategy. We define the follower’s best response as:
@@
Here, $u_{f}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ is the utility function of follower $f$. We assume that the follower breaks ties in favor of the leader.
We allow the follower to be one of $K\in\mathbb{N}_{\geq 1}$ follower types in the set $\mathscr{K}=\{f^{(1)},\dots,f^{(K)}\}$. The type of follower characterizes their utility function, $u_{f^{(i)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. We assume that the leader knows the set of possible follower types a priori, but not the realized type. The leader aims to maximize her own utility, $u:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. By slight abuse of notation, we sometimes use the shorthand
$u(\mathbf{z},\mathbf{x},a_{f})=\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f}).$
We denote a Stackelberg game with side information as a tuple, $\pazocal{G}=(\pazocal{A},\pazocal{A}_{f},\pazocal{Z},u,u_{f^{(1)}},\ldots,u_{f^{(K)}})$, consisting of the leader’s action set, the follower’s action set, the set of possible contexts, as well as the leader’s and followers’ utility functions.
@@
For the remainder of the paper and without loss of generality, we will consider $\mathscr{K}$ to be a set of distinct follower types. We call any two follower types distinct, if the optimal leader’s strategy against them is not the same for every context.

### v2

@@ Abstract
We study structured Stackelberg games, in which both players (the leader and the follower) observe contextual information about the state of the world at time of play. The leader plays against one of a finite number of followers, but the follower’s type is not known until after the game has ended. Importantly, we assume a fixed relationship between the contextual information and the follower’s type, thereby allowing the leader to leverage this additional structure when deciding her strategy. Under this setting, we find that standard learning theoretic measures of complexity do not characterize the difficulty of the leader’s learning task. Instead, we introduce a new notion of dimension, the Stackelberg-Littlestone dimension, which we show characterizes the instance-optimal regret of the leader in the online setting. Based on this, we also provide a provably optimal learning algorithm. We extend our results to the distributional setting, where we use two new notions of dimension, the $\gamma$-Stackelberg-Natarajan dimension and $\gamma$-Stackelberg-Graph dimension. We prove that these control the sample complexity lower and upper bounds respectively, and we design a simple, improper algorithm that achieves the upper bound.
@@ 1 Introduction
Many strategic interactions between a principal and an agent can be cast as Stackelberg games [38]. In a Stackelberg game, the principal (also called the leader) commits to a possibly randomized strategy, before the agent (also called the follower) observes this strategy and best-responds to it. Stackelberg games have had significant real-world impact over the past decade, particularly in security domains [37]. For example, when protecting endangered species [39], a wildlife agency determines ranger patrol locations before poachers select targets. Coast guards decide on shoreline patrol routes in advance of illegal fishers’ actions [17] and, airport security [33] chooses checkpoint screening strategies before smugglers decide which entry points to use.
@@
In many of these settings, the state of the world at time of play can greatly influence the utilities of the players and subsequently the outcome of the game. For example, patrolling some areas of the airport might be deemed of higher- or lower-importance based on the schedule of flights and the number of passengers on board. Similarly, weather patterns may make it harder to reach animals in certain parts of the park, while migration patterns might make others more vulnerable. When contextual information becomes available over time, leveraging it will generally help the leader deploy more effective strategies. This raises a natural question: given that the state of the world might evolve over time, how should the leader in a Stackelberg game design a strategy that adapts to new information coming in over time?
@@
Previous work [27, 13] has formalized such settings as repeated Stackelberg games with side information, where at each interaction both players observe contextual information that affects their utilities. In this setting, followers in different rounds may have different types that determine their payoff for each outcome of the game. Harris et al. [27] show that as long as either the sequence of followers or the sequence of contexts are drawn i.i.d. from a fixed distribution, no-regret learning is possible. That is, the leader is guaranteed to achieve nearly as much utility as the best fixed policy in hindsight. In the fully adversarial regime, however, they show that regret can grow linearly with the time horizon. Interestingly, this lower bound follows from the adversary’s ability to encode an arbitrarily complex relationship between the sequence of contexts and the follower types.
@@
Naturally, there are many cases in which the relationship between the contextual information and the follower about to appear next is not arbitrarily complex, but instead exhibits some structure. In fact, in many settings the contextual information can be predictive of the follower’s type. For example, in wildlife protection, tire tracks might signal the presence of poachers in a particular area of the park, and market demand trends might be indicative of which animals’ poachers are now more likely to appear. Similarly in airport and border security, current affairs might make it more or less likely that certain smuggler groups are active. Motivated by this observation, we study how the leader can learn effectively when there is additional structure between the contextual information and the follower types, thereby circumventing the previous impossibility result.
@@ 1.1 Our Contributions
We revisit the framework of Harris et al. [27] and take a learning theoretic approach to analyzing the leader’s task in predicting the follower type from contexts. Importantly, we find that existing notions of complexity do not capture the difficulty of learning in this setting—we obtain (tight) Stackelberg regret and sample complexity guarantees that are parameterized by new complexity measures and that directly exploit the structure of Stackelberg games. Concretely, our main contributions in this work are as follows.
Instance-optimal Stackelberg regret bounds.
@@
We initiate the study of online learning in Stackelberg games with side information, under the structural assumption that there exists a fixed, unknown mapping between contexts and follower types. In particular, we study how the leader, given access to a predefined hypothesis class, can leverage it to learn. We study the realizable setting, under which the class includes an optimal predictor of the followers’ types. Then the leader’s task is to learn an optimal policy on the fly while achieving as high cumulative utility as possible.
Naturally, the difficulty of the leader’s learning task scales with the complexity of the hypothesis class. In the learning theory literature, the (multiclass) Littlestone dimension [30, 22] is used to characterize the complexity of a class in settings with a finite set of labels (which in structured Stackelberg games corresponds to the follower types). Interestingly, we prove that the Littlestone dimension can give arbitrarily loose regret bounds (Theorem˜3.5). Our construction leverages the fact that the Littlestone dimension is completely blind to the utility space defined by the Stackelberg game, which ultimately determines the performance of the leader.
Based on these insights, we introduce a new notion of complexity, the Stackelberg-Littlestone dimension (Definitions˜3.6 and 3.7, which directly exploits the interaction between the hypothesis class and the utility space of the game. We base our definition on shattered trees, which are the workhorse of classic online learning. The challenge in our setting relies on having a finite set of follower types, yet a continuous utility function, which requires a much more nuanced approach to constructing the trees. We show that the Stackelberg-Littlestone dimension exactly characterizes the difficulty of learning in our setting (Theorem˜3.8), and in particular gives instance-optimal regret bounds. That is, for any hypothesis class and any Stackelberg game, the Stackelberg-Littlestone dimension identifies exactly the optimal regret. The proof of our upper bound Theorem˜3.9 is constructive (Algorithm˜1), which yields a complete characterization of learnability in online structured Stackelberg games.
PAC learning guarantees.
We introduce and formalize the study of distributional learning in structured Stackelberg games (Section˜4). The leader has access to a set of (context, follower type) examples, where the contexts are drawn from some fixed, unknown distribution and the follower types are determined by a true hypothesis in a given hypothesis class. Upon the draw of a new context, the leader’s goal is to deploy a strategy such as to maximize her expected payoff.
We show generalization guarantees with Probably Approximately Correct (PAC)-style bounds. We introduce the notion of $\gamma$-Stackelberg-Natarajan dimension (Definitions˜4.9 and 4.10) and $\gamma$-Stackelberg-Graph dimension (Definitions˜4.12 and 4.13), which respectively control the lower and upper bounds on the sample complexity (Theorems˜4.11 and 4.14). Our definitions are inspired by analogous notions in classification and regression [31, 3, 21], yet we give a tailored approach to what constitutes a “mistake”. This leverages the discreteness of the set of follower types to determine how much hypotheses “differ” in a minmax sense. We give a simple, improper algorithm (Algorithm˜2) that with high probability achieves (almost) optimal expected utility with as many samples as our sample complexity upper bound.
We also analyze a simpler, context-free setting, in which the training set is only over the follower types. Using tools from data-driven algorithm design, we show strong generalization guarantees Theorem˜4.5.
Optimal strategy misprediction bounds.
In the Appendix, we study structured Stackelberg games under a relaxed benchmark that seeks to minimize the total rounds in which the leader does not play the optimal strategy against the realized follower type. We give results for both the online (Appendix˜B) and distributional setting (Section˜B.2), and in both cases, we show stronger guarantees than those achieved by applying existing approaches from learning theory. These results are both conceptually and quantitatively superseded by the ones we present in the main body and we include them for completeness and further insight.
We summarize our results in  Table˜1.
Setting: Online Distributional
Dimension Definitions˜3.6 and 3.7 Definitions˜4.10 and 4.13
Lower Bound Theorem˜3.8 Theorem˜4.11
Upper Bound Theorem˜3.9 Theorem˜4.14
Algorithm Algorithm˜1 Algorithm˜2
Table 1: Summary of our main results.
@@ 1.2 Related Work
Learning in Stackelberg games. The problem of learning an optimal strategy for the leader with a single unknown follower type has been studied by Letchford et al. [29] and Blum et al. [16]. These results have been extended by Peng et al. [32], who give faster rates and almost-matching lower bounds. Recently, Bacchiocchi et al. [4] relax some structural assumptions from previous work, including tie-breaking rules for the follower, and revise worst-case sample complexity results of learning the leader’s optimal mixed strategy.
@@
Our main focus is on online learning in Stackelberg games. Online Stackelberg games with multiple follower types were introduced in Balcan et al. [7]. The authors show that when a series of unknown attacker types is chosen adversarially, playing the well-known Hedge algorithm  [25] over a carefully selected set of mixed strategies yields a no-regret algorithm.
@@
Our departing point is the work of Harris et al. [27], in which the authors study the problem of online learning in Stackelberg games with side information. The authors show that when either the sequence of follower types or the sequence of contexts are drawn independently from some fixed, unknown distribution, no-regret learning is indeed possible. Follow-up work [13] obtains better regret guarantees under bandit feedback. However, when both sequences are chosen adversarially, regret can scale linearly in the worst case, as the problem can be reduced from online learning linear thresholds. In contrast to this line of work, we circumvent this impossibility result by making an additional structural assumption about the relationship between contexts and follower types—namely that it belongs to a class of functions that the leader has access to. Under this assumption, we show that online learning is indeed possible, and our guarantees depend on an appropriate measure of complexity of this function class.
@@
Multiclass learnability. Our work can also be viewed through the lens of multiclass learnability, as it suffices for the leader to learn the hypothesis class of the function mapping the contexts to a (finite) set of follower types. Then the leader can compute the optimal strategy against the predicted follower. Multiclass prediction has been studied extensively both in the ditributional ([31, 22, 18]) and in the online setting ([22, 26]). Interestingly, we show that learning in Stackelberg games with contexts is at most as hard as multiclass prediction, and in some cases, can be much easier. For example, we give a problem instance the leader can learn to play optimally under a hypothesis class that has infinite Littlestone dimension.
Online Learning beyond Multiclass Prediction. There has been tremendous interest in understanding online learnability in many settings of interest. Notably, in online realizable regression Daskalakis and Golowich [23] designed a proper algorithm that achieved near-optimal cumulative absolute loss with respect to the sequential fat-shattering dimension of the hypothesis class. Our work is more closely related to Attias et al. [3], who introduce the online (resp. $\gamma$-Graph) dimension of a function class and show that it tightly characterizes online regression (resp. distributional learning) in the realizable setting up to a factor of 2. While our setting also considers utility functions with continuous image sets, the dimensions we introduce explicitly exploit the discrete structure of the Stackelberg game that arise from the finite set of follower types. Another recent line of work that captures incentives in online learning is online strategic classification. In this setting, agents can modify their features according to a directed graph over the feature space, before providing these features to the learner. The authors use a new combinatorial measure, the Strategic Littlestone Dimension, to achieve instance optimal mistake bounds. While seemingly similar, this measure of complexity is fundamentally distinct from the one we use to derive our results.
Learning piecewise Lipschitz functions. Given the piecewise linear structure of the leader’s utility function, our work also connects to the line of work on learning functions that are piecewise Lipschitz [8, 36, 10, 11, 9]. We draw from this literature in part of our study that explores a simpler distributional setting without contexts. The well-structured payoff function of the leader against a fixed follower type allows us to adopt tools from Balcan et al. [9] and get better generalization guarantees compared to prior work that studied a similar setting [29].
@@ 2 Problem Setting
We use $[N]:=\{1,2,\ldots,N\}$ to denote the set of natural numbers up to and including $N$. We use $\Delta(\pazocal{A})$ to denote the probability simplex over a (finite) set $\pazocal{A}$. For a given string we use inclusive slice indexing such that $x[i:j]$ are the elements of string $x$ from position $i$ to $j$. Dropping either $i$ or $j$ means that the slice extends from the beginning of the string or to the end, respectively. We use negative indexing to denote the elements of a string starting at the end i.e., $x[-1]$ is the last element of $x$.
We consider a Stackelberg game between a leader and a follower. Before the game starts, some contextual information $\mathbf{z}\in\pazocal{Z}\subseteq\mathbb{R}^{d}$ is revealed to both players. The leader plays first and commits to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$, where $\pazocal{A}$ is the set of finite actions available to the leader. The follower best-responds to the leader’s strategy and the context by playing some action $a_{f}\in\pazocal{A}_{f}$, where $\pazocal{A}_{f}$ is the finite set of follower actions. [^2] Without loss of generality the follower’s best-response will always be a pure strategy. We define the follower’s best-response as
@@
Here, $u_{f}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ is the utility function of follower $f$. In case of ties, we assume that there is a fixed and known ordering over follower actions that determines the follower’s tie-breaking rule.
We allow the follower to be one of $K\in\mathbb{N}_{\geq 1}$ follower types in the set $\{f^{(1)},\dots,f^{(K)}\}$. The type of follower characterizes their utility function, $u_{f^{(i)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. We assume that the leader knows the set of possible follower types a priori, but not the realized type and aims to maximize her own utility, $u:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. By slight abuse of notation, we sometimes use the shorthand $u(\mathbf{z},\mathbf{x},a_{f})=\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f}).$
We will denote a Stackelberg game as a tuple, $\pazocal{G}=\left(\pazocal{A},\pazocal{A}_{f},\pazocal{Z},u,u_{f^{(1)}},\ldots,u_{f^{(K)}}\right)$, consisting of the leader action set, the follower action set, the set of possible contexts, as well as the leader and followers’ utility functions.
@@
For the remainder of the paper and without loss of generality, we will consider $[K]$ to be a set of distinct follower types. We call any two follower types distinct, if the optimal leader’s strategy against them is not the same for every context.